# Rethinking MUSHRA: Addressing Modern Challenges in Text-to-Speech Evaluation

### v2

@@ Abstract
Despite rapid advancements in TTS models, a consistent and robust human evaluation framework is still lacking. For example, MOS tests fail to differentiate between similar models, and CMOS’s pairwise comparisons are time-intensive. The MUSHRA test is a promising alternative for evaluating multiple TTS systems simultaneously, but in this work we show that its reliance on matching human reference speech unduly penalises the scores of modern TTS systems that can exceed human speech quality. More specifically, we conduct a comprehensive assessment of the MUSHRA test, focusing on its sensitivity to factors such as rater variability, listener fatigue, and reference bias. Based on our extensive evaluation involving 492 human listeners across Hindi and Tamil we identify two primary shortcomings: (i) reference-matching bias, where raters are unduly influenced by the human reference, and (ii) judgement ambiguity, arising from a lack of clear fine-grained guidelines. To address these issues, we propose two refined variants of the MUSHRA test. The first variant enables fairer ratings for synthesized samples that surpass human reference quality. The second variant reduces ambiguity, as indicated by the relatively lower variance across raters. By combining these approaches, we achieve both more reliable and more fine-grained assessments. We also release Mango, a massive dataset of 246,000 human ratings, the first-of-its-kind collection for Indian languages, aiding in analyzing human preferences and developing automatic metrics for evaluating TTS systems.
@@ 1 Introduction
The above problem is well studied for the Mean Opinion Scores (MOS) test (Wester et al., 2015; Finkelstein et al., 2023; Kirkland et al., 2023; Le Maguer et al., 2024) which has received much constructive criticism over the past few years. Specifically, in a MOS test, listeners assess each system independently, which can result in an inability to accurately capture the subtle relative differences between similar systems. This poses a significant challenge in modern TTS evaluation where systems that perform equally well need to be compared against each other. To address these issues some of the recent works rely on CMOS tests (Loizou, 2011). However, this test is costly and time-consuming as it involves ($N\choose 2$ comparisons between all pairs of $N$ systems.
@@ 3 MANGO: A Corpus of Human Ratings for Speech
We introduce a new dataset, Mango: MUSHRA Assessment corpus using Native listeners and Guidelines to understand human Opinions at scale. It is a first-of-its-kind collection for any Indian language, comprising 246,000 human ratings of TTS systems and ground-truth human speech in both Hindi and Tamil, making it an expensive endeavour but one that we hope will contribute meaningfully to research in speech evaluation. Given the shortcomings of Mean Opinion Score (MOS) and Comparative Mean Opinion Score (CMOS) tests, our goal is to critically examine a promising alternative—the MUSHRA test—by conducting a large-scale evaluation involving multiple raters, systems, and languages. To do so, we adopt the standard MUSHRA test (ITU-R, 2015). Raters evaluate multiple stimuli on each page, including an explicit (mentioned) reference that serves as a benchmark for high-quality speech, along with an anchor and implicit (hidden) reference to calibrate judgments. Each stimulus is rated on a continuous scale from 0 to 100, which is also discretized: 100-80 (Excellent), 80-60 (Good), 60-40 (Fair), 40-20 (Poor), and 20-0 (Bad). We describe our evaluation setup below, and provide the detailed instructions provided to participants in Appendix A.1.
@@ 3.1 Online Annotation Platform
We enhance the webMUSHRA (Schoeffler et al., 2018) platform to address its key limitations. Specifically, we modify a fork (Pauwels et al., 2021) and introduce session management to enable saving test progress and thereby allowing for breaks for listeners. We integrate consent forms, and controls, such as ensuring raters listen to all audio samples in their entirety, to ensure more reliable ratings. We also integrate an event-tracking system to analyze time spent per page.
@@ 3.2 Synthesizing Speech Samples for Annotation
To generate samples for TTS evaluation, we train TTS systems on the Hindi and Tamil subsets of the IndicTTS database (Baby et al., 2016). Each language consists of recordings from a female and male speaker (Hindi: 20.17 hours; Tamil: 20.59 hours). We train FastSpeech2 (FS2) (Ren et al., 2021) with HiFiGAN v1 (Kong et al., 2020) and VITS (Kim et al., 2021) from scratch on the train-test splits using hyper-parameters suggested in a recent study (Kumar et al., 2023). We finetune StyleTTS2 (ST2) (Li et al., 2023) from the LibriTTS checkpoint.
@@ 3.3 Annotation Process and Dataset Statistics
To ensure reliable evaluation, we recruited native speakers of the target languages through reputable recruitment agencies. These agencies played a vital role in guaranteeing participant demographics aligned with the target language of each test. Please refer to Section A.7 for details on recruitment, consent and compensation. Once recruited, the annotators underwent a comprehensive training process comprising multiple sessions aimed at familiarizing them with the evaluation platform, test interface, and evaluation criteria. Guidelines were clearly explained, and any doubts were addressed to ensure that all participants had a uniform understanding of the evaluation process. Additionally, a structured review process was implemented, wherein participants initially rated five pilot samples. This phase allowed them to seek clarifications, provide feedback, and ensure their understanding of the evaluation criteria before proceeding to rate the 100 test samples. This approach not only improved their confidence but also helped to standardize the assessment process across all evaluators.
With the above process we collected 246,000 human ratings for TTS systems. Table 1 shows the demographic distribution, the number of participants and the overall number of ratings across all MUSHRA tests and our proposed variants (MUSHRA-NMR, MUSHRA-DG, MUSHRA-DG-NMR) which are described in Section 5.
@@ 4.1 Is MUSHRA A Reliable Test?
In Table 3, we present the results of the MUSHRA test among 3 systems and find VITS and ST2 score highest in Hindi and Tamil respectively. Surprisingly, all systems attain scores in the “Good” bin with MUSHRA scores between 60 and 80, while the reference surpasses all systems with scores in the “Excellent" bin. Given that state-of-the-art TTS systems are able to reach quality on par with references, one would expect a much smaller gap between the reference and systems. To confirm this, we conduct the more reliable but expensive CMOS test with 15 listeners in each language. In this test, we ask the rater to compare a given system, such as VITS, with a reference audio sample. The rater evaluates both the reference and the output from a system being tested without prior knowledge of which audio sample corresponds to which system, ensuring an unbiased comparison. The raters assign a single score ranging from -3 to +3 in increments of 0.5. A score of -3 indicates that System A is much worse than System B . A score of +3 indicates that System A is much better than System B. A score of 0 means that both systems are equal in quality. As seen from the scores in Table 3, CMOS indicates that the outputs synthesised by VITS and ST2 are very close in quality to the reference in Hindi and Tamil respectively, while MUSHRA scores do not reflect this at all. We hypothesize that listeners in the MUSHRA test are subject to various biases, one of which we term the reference-matching bias. This bias may lead to situations where systems that perform comparably to or better than the reference are rated less favorably, as listeners tend to focus on aligning their ratings with the reference outputs while evaluating the systems. While this may have been acceptable when TTS systems lagged behind human speech quality, it is undesirable in the current scenario where modern TTS systems often exceed the reference in aspects like naturalness and prosody Li et al. (2023); Shen et al. (2024). This suggests that the MUSHRA test, in its conventional form, may no longer be sufficient for evaluating state-of-the-art TTS systems. Instead, alternative methodologies, such as the variants we propose in Section 5, may help ensure more fair and accurate assessments.
Figure 1: Visualization of the MUSHRA score distributions per rater across three systems— FS2, ST2, and VITS, along with the reference (REF) and anchor (ANC) for Hindi. Each boxplot represents ratings (0-100) across all test utterances for a system by one rater. The substantial heights of some boxplots indicate significant variance in the scores of that rater. The variation in the means of the boxplot across raters suggests a high level of inter-rater variance. Raters are sorted in ascending order of their mean scores for the reference.
@@ 4.2 How reliable is the mean statistic in MUSHRA scores?
As mentioned earlier, each rater rates 100 utterances. In Figure 1, we use box-plots to visualize the distribution of MUSHRA scores (y-axis) for each rater (x-axis) across these utterances for each system, including the reference and anchor. While we acknowledge that the figure may appear overwhelming, we believe it is crucial for conveying the comprehensive view across both raters and utterances. We make two important observations from the figure. First, the individual box-plots have a high variance indicating that the same rater rates the system very differently across utterances. Second, looking at the means of the box-plots across different raters, we observe that there is a high variance in the means, indicating ambiguity in the perception of the MUSHRA labels across raters. We refer to this phenomenon as judgement ambiguity. This highlights the shortcomings of reporting mean statistics for MUSHRA scores, even when reported with confidence intervals (CI).
@@ 4.4 What is the impact of rejecting raters per standard MUSHRA protocol?
Traditionally, MUSHRA employs a rater rejection criterion, wherein raters scoring the hidden reference (HR) below a threshold ($\lambda$) more than 15% of the time are rejected. This rejection rule stems from the inherent assumption that the HR is the gold standard, which is not true in modern TTS settings where TTS systems (Li et al., 2023; Shen et al., 2024) that is either on par with or surpasses the reference. In such cases, a rater consistently scoring the HR lower might not necessarily reflect unreliability, but rather a nuanced perception of the reference’s limitations compared to the evaluated systems. This is reinforced by observations from Table 3 where raters clearly prefer ST2 over reference for Tamil with a mean CMOS score of 0.24. This observation is further supported by the MUSHRA scores shown in Figure 3. The table shows that while rejecting raters based on the conventional MUSHRA criterion does not affect system rankings, it does notably shift scores. Specifically, system scores decrease with increasing $\lambda$, while reference scores increase. This trend hints at the reference-matching bias wherein raters who give the HR high scores might be unconsciously matching system samples to the mentioned reference, rather than rating based on their absolute perception of quality.
@@ 4.5 How does Anchor affect scores?
In MUSHRA tests, the anchor serves the purpose of setting the expectation of what a Fair sample sounds like. Typically, the anchor is created by minimally degrading the ground-truth by first downsampling it to 3.5 kHz and then upsampling it to 24 kHz. We refer to such an anchor as Anchor-X. In Table 3, the mean scores for Anchor-X in Hindi indicate that this anchor performs significantly better than all other systems, attaining a high score of 70.81. We believe these scores are explained by the resampling strategy used to create the Anchor-X, which introduces some artifacts in the audio but retains similar naturalness to the reference, especially in terms of prosody. Once again, this intuition indicates the tendency of raters to rate systems that match the reference with higher scores (reference-matching bias). We conclude that using this anchor may not be ideal, as it can lead to potentially “Excellent” TTS systems being unfairly rated as “Good”. Essentially, the raters may perceive that if one of systems (anchor, in this case) sounds very similar to the reference, there is little justification for rating other systems highly.

### v3

@@ Abstract
Despite rapid advancements in TTS models, a consistent and robust human evaluation framework is still lacking. For example, MOS tests fail to differentiate between similar models, and CMOS’s pairwise comparisons are time-intensive. The MUSHRA test is a promising alternative for evaluating multiple TTS systems simultaneously, but in this work we show that its reliance on matching human reference speech unduly penalises the scores of modern TTS systems that can exceed human speech quality. More specifically, we conduct a comprehensive assessment of the MUSHRA test, focusing on its sensitivity to factors such as rater variability, listener fatigue, and reference bias. Based on our extensive evaluation involving 492 human listeners across Hindi and Tamil we identify two primary shortcomings: (i) reference-matching bias, where raters are unduly influenced by the human reference, and (ii) judgement ambiguity, arising from a lack of clear fine-grained guidelines. To address these issues, we propose two refined variants of the MUSHRA test. The first variant enables fairer ratings for synthesized samples that surpass human reference quality. The second variant reduces ambiguity, as indicated by the relatively lower variance across raters. By combining these approaches, we achieve both more reliable and more fine-grained assessments. We also release Mango [^1] The dataset is publicly available at https://huggingface.co/datasets/ai4bharat/MANGO., a massive dataset of 246,000 human ratings, the first-of-its-kind collection for Indian languages, aiding in analyzing human preferences and developing automatic metrics for evaluating TTS systems .
@@ 1 Introduction
The above problem is well studied for the Mean Opinion Scores (MOS) test (Wester et al., 2015; Finkelstein et al., 2023; Kirkland et al., 2023; Le Maguer et al., 2024) which has received much constructive criticism over the past few years. Specifically, in a MOS test, listeners assess each system independently, which can result in an inability to accurately capture the subtle relative differences between similar systems. This poses a significant challenge in modern TTS evaluation where systems that perform equally well need to be compared against each other. To address these issues some of the recent works rely on CMOS tests (Loizou, 2011). However, this test is costly and time-consuming as it involves $\binom{N}{2}$ comparisons between all pairs of $N$ systems.
@@ 2 Related Work
Perceptual Judgments. TTS evaluation primarily relies on naturalness, which, while lacking a universally agreed-upon definition (Shirali-Shahreza & Penn, 2023), generally refers to how closely synthesized speech resembles human speech in perception. This concept aligns with standard evaluation practices and is typically assessed through subjective listening tests such as MOS (Łańcucki, 2021; Ren et al., 2021; Kim et al., 2021), CMOS (Shen et al., 2024; Tan et al., 2024; Li et al., 2023) or MUSHRA (Lajszczak et al., 2024), where listeners rate speech samples based on their perceived realism and overall fidelity.
@@ 3 MANGO: A Corpus of Human Ratings for Speech
We introduce a new dataset, Mango: MUSHRA Assessment corpus using Native listeners and Guidelines to understand human Opinions at scale. It is a first-of-its-kind collection for any Indian language, comprising 246,000 human ratings of TTS systems and ground-truth human speech in both Hindi and Tamil, making it an expensive endeavour but one that we hope will contribute meaningfully to research in speech evaluation. Given the shortcomings of Mean Opinion Score (MOS) and Comparative Mean Opinion Score (CMOS) tests, our goal is to critically examine a promising alternative—the MUSHRA test—by conducting a large-scale evaluation involving multiple raters, systems, and languages. To do so, we adopt the standard MUSHRA test (ITU-R, 2015).
Raters evaluate multiple stimuli on each page, including an explicit (mentioned) reference that serves as a benchmark for high-quality speech, along with an anchor and implicit (hidden) reference to calibrate judgments. Each stimulus is rated on a continuous scale from 0 to 100, which is also discretized: 100-80 (Excellent), 80-60 (Good), 60-40 (Fair), 40-20 (Poor), and 20-0 (Bad). We describe our evaluation setup below, and provide the detailed instructions provided to participants in Appendix A.1.
3.1 MUSHRA-Based Annotation Framework
To evaluate TTS systems, we use the standard MUSHRA test ITU-R (2015), a methodology originally developed for codec evaluation and widely used in speech and audio quality assessment. MUSHRA is particularly suited for listeners with normal hearing and experience in critical listening, as it requires evaluating multiple stimuli side by side while referencing a high-quality benchmark. Raters assess multiple stimuli per page, including an explicit reference (a high-quality speech benchmark), an anchor (a lower-quality version for calibration), and an implicit (hidden) reference to normalize judgments. Each stimulus is rated on a continuous 0–100 scale, discretized into five categories: 100-80 (Excellent), 80-60 (Good), 60-40 (Fair), 40-20 (Poor), and 20-0 (Bad). To ensure reliable and consistent evaluations, MUSHRA follows established guidelines according to the standard. Listeners must complete pre-test training to familiarize themselves with impairments and test signals. Stimuli should be around 10 seconds long, with a maximum of 12 seconds, to minimize listener fatigue and enhance response stability. An assessor’s responses are excluded if they rate the hidden reference below 90 in more than 15% of test items. Tests must use either headphones or loudspeakers, but not both within a session, ensuring consistency across participants. When listening conditions are technically and behaviorally controlled, data from as few as 20 subjects can yield statistically reliable conclusions. These measures help maintain the rigor and reproducibility of MUSHRA-based speech evaluations, making it a robust choice for large-scale TTS assessment.
Comparison with Alternative Tests. MUSHRA offers several advantages over Mean Opinion Score (MOS) and Comparative MOS (CMOS) for speech quality assessment. MOS, despite its widespread use, has been criticized for its poor reliability and inability to distinguish between similar-sounding systems due to its single-stimulus nature and the absence of an explicit reference. While CMOS provides pairwise comparisons, it is inherently limited to evaluating only two systems at a time, making large-scale assessments both expensive and difficult to scale. In contrast, MUSHRA enables simultaneous comparisons across multiple systems, allowing listeners to switch between samples and directly compare quality, which enhances sensitivity to subtle differences that MOS often fails to capture. The continuous 0–100 scale provides finer granularity than MOS’s coarse 5-point rating, allowing for more nuanced judgments. Additionally, MUSHRA is more efficient than exhaustive pairwise CMOS tests, as it allows multiple systems to be evaluated in a single trial rather than requiring numerous pairwise comparisons. Another key strength is reference-based calibration, where listeners rate systems relative to a known high-quality reference and an anchor, with the aim of reducing scoring variability. Ribeiro et al. (2015) further confirm MUSHRA’s advantages, demonstrating that it more effectively distinguishes between models compared to MOS, reinforcing its suitability for evaluating modern TTS systems.
Evolution of MUSHRA. Over time, researchers have refined the MUSHRA framework to better align with the evolving demands of speech synthesis evaluation. While the standard MUSHRA test remains widely used, several variations have emerged, modifying key aspects such as reference anchoring and test structure to address specific challenges in assessing TTS quality. For instance, BaseTTS (Lajszczak et al., 2024) follows the MUSHRA framework but omits anchors, suggesting that explicit anchoring may not always be necessary for reliable judgments. NaturalSpeech2 (Shen et al., 2024) further questions the necessity of strict reference matching by employing a MUSHRA-like CMOS setup without an explicit reference. Additionally, Expressive-MUSHRA has been introduced to evaluate expressiveness in speech synthesis (Huybrechts et al., 2021; Varadhan et al., 2024), highlighting MUSHRA’s flexibility in capturing prosodic variation and expressive nuances. Beyond these structural modifications, studies have also examined MUSHRA’s effectiveness and limitations. For example, Merritt et al. (2018) critique the traditional 3.5 kHz anchor, arguing that it may not be ideal for modern speech synthesis, as distortions in neural TTS systems differ from those originally considered in codec evaluations. These adaptations reflect a growing recognition that MUSHRA, while valuable, may require refinements to remain effective for contemporary TTS evaluation. Recognizing this need, we systematically explore alternative test designs and their impact on evaluation reliability.
@@ 3.2 Online Annotation Platform
We enhance the webMUSHRA (Schoeffler et al., 2018) platform to address its key limitations. Specifically, we modify a fork (Pauwels et al., 2021) and introduce session management to enable saving test progress, and thereby allowing for breaks for listeners. We integrate consent forms and controls, such as ensuring raters listen to all audio samples in their entirety, to ensure more reliable ratings. We also integrate an event-tracking system to analyze the time spent per page.
@@ 3.3 Synthesizing Speech Samples for Annotation
To generate samples for TTS evaluation, we train TTS systems on the Hindi and Tamil subsets of the IndicTTS database (Baby et al., 2016). Each language consists of recordings from a female and male speaker (Hindi: 20.17 hours; Tamil: 20.59 hours). We train FastSpeech2 (FS2) (Ren et al., 2021) with HiFiGAN v1 (Kong et al., 2020) and VITS (Kim et al., 2021) from scratch on the train-test splits using hyper-parameters suggested in a recent study (Kumar et al., 2023b). We finetune StyleTTS2 (ST2) (Li et al., 2023) from the LibriTTS checkpoint.
@@ 3.4 Annotation Process and Dataset Statistics
To ensure reliable evaluation, we recruited native speakers of the target languages through reputable recruitment agencies. These agencies played a vital role in guaranteeing participant demographics aligned with the target language of each test. Please refer to Section A.9 for details on recruitment, consent, and compensation. Once recruited, the annotators underwent a comprehensive training process comprising multiple sessions aimed at familiarizing them with the evaluation platform, test interface, and evaluation criteria. As part of this process, demo calls were conducted, where participants were introduced to the rating interface and guided through a sample rating task. These sessions allowed participants to ask questions and receive clarifications in real time. Additionally, a structured review process was implemented, wherein participants initially rated five pilot samples. This phase allowed them to seek clarifications, provide feedback, and ensure their understanding of the evaluation criteria before proceeding to rate the 100 test samples. All instructions provided to participants are included in Appendix A.6 for transparency. This approach not only improved their confidence, but also helped standardize the assessment process across all evaluators.
With the above process, we collected 246,000 human ratings for TTS systems. We also collect demographic information from a majority of the 492 participants, while a subset of 21 individuals preferred not to disclose their details. Table 1 shows the demographic distribution, the number of participants, and the total number of ratings across all MUSHRA tests and our proposed variants (MUSHRA-NMR, MUSHRA-DG, MUSHRA-DG-NMR) which are described in Section 5. The table reflects responses from participants who disclosed demographic details, while a small subset preferred not to share this information.
@@ 4.1 Is MUSHRA A Reliable Test?
In Table 3, we present the results of the MUSHRA test among 3 systems and find VITS and ST2 score highest in Hindi and Tamil respectively. Surprisingly, all systems attain scores in the “Good” bin with MUSHRA scores between 60 and 80, while the reference surpasses all systems with scores in the “Excellent" bin. Given that state-of-the-art TTS systems are able to reach quality on par with references, one would expect a much smaller gap between the reference and systems. To confirm this, we conduct the more reliable but expensive CMOS test with 15 listeners in each language. In this test, we ask the rater to compare a given system, such as VITS, with a reference audio sample. The rater evaluates both the reference and the output from a system being tested without prior knowledge of which audio sample corresponds to which system, ensuring an unbiased comparison. The raters assign a single score ranging from -3 to +3 in increments of 0.5. A score of -3 indicates that System A is much worse than System B . A score of +3 indicates that System A is much better than System B. A score of 0 means that both systems are equal in quality. As seen from the scores in Table 3, CMOS indicates that the outputs synthesised by VITS and ST2 are very close in quality to the reference in Hindi and Tamil respectively, while MUSHRA scores do not reflect this at all. We hypothesize that listeners in the MUSHRA test are subject to various biases, one of which we term the reference-matching bias. This bias may lead to situations where systems that perform comparably to or better than the reference are rated less favorably, as listeners tend to focus on aligning their ratings with the reference outputs while evaluating the systems. While this may have been acceptable when TTS systems lagged behind human speech quality, it is undesirable in the current scenario where modern TTS systems often exceed the reference in aspects like naturalness and prosody (Li et al., 2023; Shen et al., 2024). This suggests that the MUSHRA test, in its conventional form, may no longer be sufficient for evaluating state-of-the-art TTS systems. Instead, alternative methodologies, such as the variants we propose in Section 5, may help ensure more fair and accurate assessments.
Further evidence questioning the role of the reference as the highest standard in MUSHRA evaluations is provided in Appendix A.2, where we analyze CMOS-derived preference ratings and highlight cases where TTS systems are notably preferred over the human reference.
Figure 1: Visualization of the MUSHRA score distributions per rater across three systems— FS2, ST2, and VITS, along with the reference (REF) and anchor (ANC) for Hindi. Each boxplot represents ratings (0-100) across all test utterances for a system by one rater.While some boxplots exhibit relatively low variance, several display substantial height, indicating that certain raters assign widely varying scores to the same system across utterances. The variation in the means of the boxplot across raters suggests a high level of inter-rater variance. Raters are sorted in ascending order of their mean scores for the reference.
@@ 4.2 How reliable is the mean statistic in MUSHRA scores?
As mentioned earlier, each rater rates 100 utterances. In Figure 1, we use box-plots to visualize the distribution of MUSHRA scores (y-axis) for each rater (x-axis) across these utterances for each system, including the reference and anchor. While we acknowledge that the figure may appear overwhelming, we believe it is crucial for conveying the comprehensive view across both raters and utterances. We make two important observations from the figure. First, many the individual box-plots (especially of a system) have a high variance indicating that the same rater rates the system very differently across utterances. Second, looking at the means of the box-plots across different raters, we observe that there is a high variance in the means, indicating ambiguity in the perception of the MUSHRA labels across raters. We refer to this phenomenon as judgement ambiguity. This highlights the shortcomings of reporting mean statistics for MUSHRA scores, even when reported with confidence intervals (CI).
@@ 4.4 What is the impact of rejecting raters per standard MUSHRA protocol?
Traditionally, MUSHRA employs a rater rejection criterion, wherein raters scoring the hidden reference (HR) below a threshold ($\lambda$) more than 15% of the time are rejected. This rejection rule stems from the inherent assumption that the HR is the gold standard, which is not true in modern TTS settings where TTS systems (Li et al., 2023; Shen et al., 2024) are either on par with or surpass the reference. In such cases, a rater consistently scoring the HR lower might not necessarily reflect unreliability, but rather a nuanced perception of the reference’s limitations compared to the evaluated systems. This is reinforced by observations from Table 3 where raters clearly prefer ST2 over reference for Tamil with a mean CMOS score of 0.24. This observation is further supported by the MUSHRA scores shown in Figure 3. The table shows that while rejecting raters based on the conventional MUSHRA criterion does not affect system rankings, it does notably shift scores. Specifically, system scores decrease with increasing $\lambda$, while reference scores increase. This trend hints at the reference-matching bias wherein raters who give the HR high scores might be unconsciously matching system samples to the mentioned reference, rather than rating based on their absolute perception of quality.
@@ 4.5 How does Anchor affect scores?
In MUSHRA tests, the anchor serves the purpose of setting the expectation of what a “Fair” sample sounds like. Typically, the anchor is created by minimally degrading the ground-truth by first downsampling it to 3.5 kHz and then upsampling it to 24 kHz. We refer to such an anchor as Anchor-X. In Table 3, the mean scores for Anchor-X in Hindi indicate that this anchor performs significantly better than all other systems, attaining a high score of 70.81. We believe these scores are explained by the resampling strategy used to create the Anchor-X, which introduces some artifacts in the audio but retains similar naturalness to the reference, especially in terms of prosody. Once again, this intuition indicates the tendency of raters to rate systems that match the reference with higher scores (reference-matching bias). We conclude that using this anchor may not be ideal, as it can lead to potentially “Excellent” TTS systems being unfairly rated as “Good”. Essentially, the raters may perceive that if one of systems (anchor, in this case) sounds very similar to the reference, there is little justification for rating other systems highly.